# Detecting State Changes in Functional Neuronal Connectivity using Factorial Switching Linear Dynamical Systems

### v1

@@ 3.3 Inferring latent states using variational inference
In order to effectively learn the latent states $\Theta=\{h_{t},z_{t}\}_{t=1}^{T}$ we develop an AEVB algorithm inspired by [2]. Our algorithm makes use of a variational distribution $q(\Theta)=\prod_{t}q(h_{t})q(z_{t})$ to approximate the posterior of the model defined in Equation LABEL:eqn:our_model. We parametrize $q(\Theta)$ using a combination of neural networks,
$q(h_{t}^{(k)})=$ $\text{Concrete}\left(f_{q}\left(h_{t-1}^{(k)},y_{t}\right)\right)$ (5)
$q(z_{t}^{(k)})=$ $\text{Normal}\left(f_{z}\left(z_{t-1}^{(k)},y_{t}\right),\sigma_{q}^{2}\right).$
Here, our variational distribution $q\left(h_{t}^{(k)}\right)$ for the discrete latent states is parametrized by an RNN $f_{q}$, and the variational distribution $q\left(z_{t}^{(k)}\right)$ for the continuous latent states is parameterized by an RNN $f_{z}$. Each RNN depends both on the previous states and on the current observation.
We then optimize the ELBO, following [12]. We treat $\sigma_{p}$, $\sigma_{q}$, and $A$ as parameters to be optimized using stochastic gradient descent. To encourage sparsity in the inferred latent structure, we add an L1 penalty $\lambda\sum_{t}\sum_{k}h_{t}^{(k)}$ on the learned $h_{t}^{(k)}$ to the ELBO.
@@ 4 Experimental evaluation
In this section, we evaluate FSLDS on a series of real and synthetic datasets. For all experiments, we represent $f_{q}$, $f_{p}$ and $f_{z}$ using recurrent neural networks with 64 hidden dimensions. The temperature for the Concrete distribution was gradually annealed from 1 at a rate of 0.00003 until it reached a final value of 0.1. In the simulated data experiments, we set a learning rate of 0.01 for all parameters except for $\theta$, which initially had a learning rate of 0.1 that was reduced to 0.01 after the first 200 epochs. For real data experiments, we used a learning rate of 0.001 for all parameters, again with $\theta$ starting at 0.1 and decreasing to 0.001 after 200 epochs. We augment the ELBO with an L1 penalization term $\lambda\sum_{t}\sum_{k}h_{t}^{(k)}$, with $\lambda=0.2$; this value was found to give good performance on synthetic data. For each value of $K$, we use 75 different random initializations and select the run with the lowest ELBO.
@@
(a) Latent features $\theta_{0:4}$.
@@
(e) Combined rate $\theta_{0}e^{z_{t}^{(0)}}+\sum_{k=1}^{4}\theta_{k}h_{t}^{(k)}e^{z_{t}^{(k)}}$.
@@ 4.1 Simulated Data
We begin by exploring the performance of FSLDS in a synthetic example where the ground truth features are known (but not generated from the model). This allows us to explore the ability to recover latent structure and showcase the benefits of using a factorial model over a simpler SLDS. We generate $T=1000$ time steps, with $K=4$ latent, 16-dimensional features $\theta_{k}$, as shown in Figure 3(a). Each figure has a latent continuous trajectory $z_{t}^{(k)}$ drawn either from a straight line or a sinusoid, as shown in Figure 3(c). These were combined with manually generated binary trajectories $h_{t}^{(k)}$(Figure 3(b), Figure 3(d)) to obtain the rate shown in Figure 3(e). This rate was used to sample random Poisson observations.
In Figure 4, we show the posterior means obtained via the inference procedure described in Section 3.3. The results indicate that the Poisson rates are accurately recovered and are appropriately decomposed into the true generating features, demonstrating the model’s ability to capture the underlying structure of the data. Even though we set $K=6$, the model correctly identifies the 4 latent features shown in Figure 3. We note that, while the ConcreteFHMM uses a continuous-valued indicator $h_{t}^{(k)}$, in practice the values obtained are very close to zero or one (Figure 4(b)); we see similar behavior in all other experiments in this section.
@@
Figure 5: Latent states obtained using rSLDS. Left hand plot shows the ground truth latent states (corresponding to the twelve unique combinations of four additive features shown in Figure 3(b)). Remaining plots show the results with varying numbers of latent states, from four to sixteen.
To attempt to learn this true, non-additive structure, we run recurrent SLDS (rSLDS), using the ssm package associated with the original paper [16]. We test rSLDS with $K$= 4, 7, 10, 13 and 16 features (with the correct number being 13), taking the best ELBO out of ten random seeds. The resulting state assignments shown in Figure 5. We see that we cannot recover the ground truth. We find that rSLDS tends to miss low-occupancy features and combine multiple features. Even though most of the plots show between two and four main features, these features are not consistent between plots. The results also tend to be noisier than those obtained using FSLDS, particularly as the number of features increases.
@@ 4.2 Evaluation of microelectrode array spike train data
To evaluate the performance of our algorithm on real-world neuronal connectivity data, we look at two datasets obtained via MEA recordings. For both datasets, we use preprocessed data from [28]. To obtain this data, spikes (action potentials) were detected from the raw MEA voltage time series, based on the action potential waveform using the continuous wavelet transform method implemented in MEA-NAP, a MATLAB pipeline for MEA network analysis[28]. [^2] https://github.com/SAND-Lab/MEA-NAP The voltage timeseries were acquired at a sampling rate of 12500 Hz (human iPSC-dervied neuronal cultures) and 25000 Hz (murine hippocampal cultures). Action potentials are approximatedly 1ms in duration. For our analysis, we aggregated the spike train data to 1 second time bins.
@@ 4.2.1 Exploring properties of the inferred structure using murine hippocampal recordings
(d) Latent subgraphs $\theta_{k}$.
@@
Figure 6: Murine hippocampal data (recording TC043 DIV14). Figure 6(a) shows the spike counts (color bar, truncated to the 95th percentile value) from a 10-minute MEA recording of a days-in-vitro (DIV) 14 primary hippocampal culture. The activity observed in each electrode (columns) is shown over time (y-axis unit is seconds). Spike count binned by 1 second intervals. Figure 6(b) shows the binary feature indicators $h_{t}^{(k)}$, indicating which of the latent features are active at each time point. Figure 6(c) shows the expression level of each of the latent features $\theta_{k}$. Figure 6(d) shows the relative contribution of each electrode, in the spatial arrangement of the MEA, for a subset of the latent states $\theta_{k}$, including the background (subnetwork 0) and most active subnetworks (1-4). The relative contribution (node color) is scaled for each subnetwork.
(a) Binned spike counts.
(b) Feature indicators $h_{t}^{(k)}.$
(c) Combined activities $e^{z_{t}^{(0)}},h_{t}^{(k)}e^{z_{t}^{(k)}}$.
(d) Latent subgraphs $\theta_{k}$.
Figure 7: Murine hippocampal data (recording TC165 DIV14). Figure 7(a) shows the spike counts (color bar, truncated to the 95th percentile value) from a 10-minute MEA recording of a days-in-vitro (DIV) 14 primary hippocampal culture. The activity observed in each electrode (columns) is shown over time (y-axis unit is seconds). Spike count binned by 1 second intervals. Figure 7(b) shows the binary feature indicators $h_{t}^{(k)}$, indicating which of the latent features are active at each time point. Figure 7(c) shows the expression level of each of the latent features $\theta_{k}$. Figure 7(d) shows the relative contribution of each electrode, in the spatial arrangement of the MEA, for a subset of the latent states $\theta_{k}$, including the background (subnetwork 0) and most active subnetworks (1-4). The relative contribution (node color) is scaled for each subnetwork.
@@
We first consider qualitative analyses based on MEA recordings from two 2D murine hippocampal cultures, first published in [26], using the spike train data from [28]. We visualize the neuronal activity from these two recordings in Figures 6(a) and 7(a). The remaining subplots in Figures 6 and 7 show the latent structure obtained using FSLDS. In each case, we see the activity decomposed into distinct subgraphs (Figures 6(d) and 7(d)). Some features are active throughout most of the recording, but with varying levels of activity (e.g., features 1 and 2 in Figure 6). Others are active in short bursts, such as features 3 and 4 in Figure 6.
One concern might be that the number of inferred latent features will grow with sequence length regardless of the true latent dimensionality. To demonstrate this is not the case, in Figure 8, we concatenate the recording shown in Figure 6 with itself and re-run our analysis. We find there are still 4 main features, and the four features are almost the same as the features from using one single copy of the TC043 DIV14 recording (c.f. Figures 6(d) and 8(d)). Additionally, the activation patterns in the concatenated part is almost the same as the first half, indicating that the model does not create arbitrary patterns due to longer recordings.
(a) Binned spike counts.
(b) Feature indicators $h_{t}^{(k)}.$
(c) Combined activities $e^{z_{t}^{(0)}},h_{t}^{(k)}e^{z_{t}^{(k)}}$.
(d) Latent subgraphs $\theta_{k}$.
Figure 8: Murine hippocampal data (recording TC043 DIV14), duplicated, and concatenated, creating a 12-minute data series. Figure 8(a) shows the spike counts (truncated to the 95th percentile value) for the recording (1st copy 0-600 seconds, 2nd copy 600-1200 seconds). Figure 8(b) shows the binary feature indicators $h_{t}^{(k)}$, indicating which of the latent features are active at each time point. Figure 8(c) shows the expression level of each of the latent features $\theta_{k}$. Figure 8(d) shows the relative contribution of each electrode, in the spatial arrangement of the MEA, for a subset of the latent states $\theta_{k}$, including the background (subnetwork 0) and most active subnetworks (1-4).
(a) Binned spike counts.
(b) Feature indicators $h_{t}^{(k)}.$
(c) Combined activities $e^{z_{t}^{(0)}},h_{t}^{(k)}e^{z_{t}^{(k)}}$.
(d) Latent subgraphs $\theta_{k}$.
Figure 9: Murine hippocampal data (recordings TC043 DIV14 and TC165 DIV14) concatenated, creating a 12-minute time series. Figure 9(a) shows the spike counts (truncated to the 95th percentile value) for the two recordings (1st culture 0-600 seconds, 2nd culture 600-1200 seconds). Figures 9(b)9(c) show inferred latent structure. Figure 9(d) shows subgraphs, in the spatial arrangement of the MEA, for the subnetworks (0-6).
In contrast, Figure 9 shows the same recording concatenated with a second recording (TC165 DIV14, as shown in Figure 7). Here, we would expect the structure for the first recording to be similar to that obtained in Figure 6, but unlike the previous example, we do not expect these features to persist into the second recording. Instead, we would expect to see latent structure similar to Figure 7. We see both halves of the concatenation use completely different features—despite the fact that some electrodes, such as Electrode 36, appear relatively active in both Figures 6(a) and 7(a). We also see that the activation features remain similar to the single recordings.
@@ 4.2.2 Exploring evolution of connectivity over longer time periods
Figure 10: Number of activated features across weeks in culture. Violin plots with scatter plots of number of active features for each culture (circles or stars, n=15 cultures) at four developmental time points days-in-vitro (DIV) 14, 21, 28, and 35. Lines show developmental trajectory of number of latent features within individual cultures. Three examples are highlighted with stars and thicker colored lines: P1B2 (pink, see also Figure 12), P5A2 (orange, see also Figure 13), and P3B1 (purple, see also Figure 14).
Next, we consider a dataset of MEA recordings from human induced pluripotent stem cell (iPSC)-derived neurogenin-2 (NGN2) cortical neurons cultured on 64-electrode MEAs (n=15 cultures), first analyzed in [28]. Each culture was recorded for a period of ten minutes once per week, over a period of four weeks, at days-in-vitro (DIV)14, 21, 28 and 35 for $n=15$ cortical cultures. Spike counts were binned into one-second intervals. To investigate subnetwork development in vitro, we concatenated the data from all four weeks for each culture.
@@
In Figure 10, we show the average number of active features (or subnetworks) for each week, for each recording. In almost all cultures, we see an increase in the number of active features across the four weeks, suggesting a developmental process. This coincides with an increase in the number of active electrodes and the number of action potentials detected by these electrodes [28]. However, this pattern diverges at DIV 35, where a bimodal behaviour in the dataset is observed. While a few cultures continue to show a similar number (or increase) in the number of subgraphs at DIV 35 compared to DIV 28 (e.g., Figures 13), in other cultures, this number decreases (e.g., Figure 12). These likely represent more mature cultures in which the increase in the degree of connectivity has led to a consolidation of subgraphs between DIV 28-35 (see [28]).
@@
(d) Latent subgraphs $\theta_{k}$.
(e) Mean network activity.
@@
Figure 12: Development of human iPSC-derived NGN2 cortical neurons in culture P1B2. Figure 12(a) shows the raster plot from the concatenated spike train data from days-in-vitro (DIV) 14, 21, 28, and 35 with the spike count (binned in 1 second intervals) for the 64 electrodes (rows) over the four 10-minute recordings (x-axis unit is seconds). Figure 12(b) shows the latent features (rows) for the 10-minute MEA recordings from the four developmental time points (column). Color bar represents the probability of the feature being on (yellow) or off (purple). Figure 12(c) shows the combined activity (color bar) revealing the temporal dynamics within the on and off switching of the latent states within each MEA recording. Lighter colors (pink to yellow) show dominant features in contrast to features that are off (dark blue). Figure 12(d) shows heatmaps of the relative contribution (color) of individual electrodes (circles) to the subgraphs in the spatial arrangement of the MEA for subnetworks 0-7. Figure 12(e) shows heatmaps in the spatial arrangement of the MEA of the mean spike count (per second) per electrode (color) at the four developmental time points.
(a) Binned spike counts.
(b) Feature indicators $h_{t}^{(k)}$.
(c) Combined activities $e^{z_{t}^{(0)}},h_{t}^{(k)}e^{z_{t}^{(k)}}$.
(d) Latent subgraphs $\theta_{k}$.
(e) Mean network activity.
Figure 13: Development of human iPSC-derived NGN2 cortical neurons in culture P5A2. Figure 13(a) shows the raster plot from the concatenated spike train data from days-in-vitro (DIV) 14, 21, 28, and 35 with the spike count (binned in 1 second intervals) for the 64 electrodes (rows) over the four 10-minute recordings (x-axis unit is seconds). Figure 13(b) shows the latent features (rows) for the 10-minute MEA recordings from the four developmental time points (column). Color bar represents the probability of the feature being on (yellow) or off (purple). Figure 13(c) shows the combined activity (color bar) revealing the temporal dynamics within the on and off switching of the latent states within each MEA recording. Lighter colors (pink to yellow) show dominant features in contrast to features that are off (dark blue). Figure 13(d) shows heatmaps of the relative contribution (color) of individual electrodes (circles) to the subgraphs in the spatial arrangement of the MEA for subnetworks 0-5. Figure 13(e) shows heatmaps in the spatial arrangement of the MEA of the mean spike count (per second) per electrode (color) at the four developmental time points.
(a) Binned spike counts.
(b) Feature indicators $h_{t}^{(k)}$.
(c) Combined activities $e^{z_{t}^{(0)}},h_{t}^{(k)}e^{z_{t}^{(k)}}$.
(d) Latent subgraphs $\theta_{k}$.
(e) Mean network activity.
Figure 14: Development of human iPSC-derived NGN2 cortical neurons in culture P3B1. Figure 14(a) shows the raster plot from the concatenated spike train data from days-in-vitro (DIV) 14, 21, 28, and 35 with the spike count (binned in 1 second intervals) for the 64 electrodes (rows) over the four 10-minute recordings (x-axis unit is seconds). Figure 13(b) shows the latent features (rows) for the 10-minute MEA recordings from the four developmental time points (column). Color bar represents the probability of the feature being on (yellow) or off (purple). Figure 13(c) shows the combined activity (color bar) revealing the temporal dynamics within the on and off switching of the latent states within each MEA recording. Lighter colors (pink to yellow) show dominant features in contrast to features that are off (dark blue). Figure 13(d) shows heatmaps of the relative contribution (color) of individual electrodes (circles) to the subgraphs in the spatial arrangement of the MEA for subnetworks 0-7. Figure 13(e) shows heatmaps in the spatial arrangement of the MEA of the mean spike count (per second) per electrode (color) at the four developmental time points.
@@
In the first culture, P1B2 (Figure 12), we see an initial increase in the number of subnetworks from DIV 14-28, followed by a mild consolidation of the network structure into new activity patterns at DIV 35. This illustrates the ability of the model to detect biologically relevant changes in subnetworks as the neuronal networks mature. The model detects seven different features (Figure 12(b)12(c)), and the latent subgraphs reveal different combinations of electrodes detecting activity in these subnetworks (Figure 12(d)). From DIV 14 to DIV 28, we observe an increasing number of different features that are turned on (from 1 to 3). In addition, we see one feature introduced at DIV 21 remains active during DIV 28. However at DIV 35, one completely different feature is activated. These findings are consistent with an increase in network activity (Figure 12(e)) and density, illustrated in [28], at DIV14-28 and an increase in the mean nodal participation coefficient between DIV 28-35. When considering the spatial arrangement of the subnetworks (Figure 12(d)), we also observe a shift in participating nodes. During the first three weeks, the active nodes are more likely to be located in the middle or right side of the MEA, while in DIV35, the activity has shifted to the left. This may indicate that some subnetworks may increase or decrease in dominance in the overall network activity as the cultures mature.
@@
The second culture, P5A2 (Figure 13), clearly demonstrates an increase in network complexity from DIV 14-35 (Figure 13(b)13(c)), without the consolidation behavior seen in P1B2 at DIV 35. With no active features observed at DIV14, subnetwork 0 (Figure 13(d)) mainly comes from the baseline noise. As the neuronal networks mature, more features are activated, forming more complex subnetworks (Figure 13(c)13(d)). We hypothesize that P5A2 may be following a similar developmental trajectory to P1A3, but with a slower start (as evidenced by the lack of correlated activity at DIV 14). As with P1A3, we observe the overall network structure becoming more complex as additional subnetworks are activated.

### v2

@@ 3.3 Inferring latent states using variational inference
In order to effectively learn the latent states $\Theta=\{h_{t},z_{t}\}_{t=1}^{T}$ we develop an AEVB algorithm inspired by Becker-Ehmck et al., (2019). Our algorithm makes use of a variational distribution $q(\Theta)=\prod_{t}q(h_{t})q(z_{t})$ to approximate the posterior of the model defined in Equation LABEL:eqn:our_model. We parametrize $q(\Theta)$ using a combination of neural networks,
$q(h_{t}^{(k)})=$ $\text{Concrete}\left(f_{q}\left(h_{t-1}^{(k)},y_{t}\right),\phi\right)$ (5)
$q(z_{t}^{(k)})=$ $\text{Normal}\left(f_{z}\left(z_{t-1}^{(k)},y_{t}\right)\right).$
Here, our variational distribution $q\left(h_{t}^{(k)}\right)$ for the discrete latent states is parametrized by a neural network $f_{q}$, and the variational distribution $q\left(z_{t}^{(k)}\right)$ for the continuous latent states is parameterized by a neural network $f_{z}$. Note that $f_{z}$ parametrizes both the mean and the standard deviation of the normal distribution. Each neural network depends both on the previous states and on the current observation. By default we use MLPs for $f_{q}$, $f_{p}$ and $f_{z}$, as it is the simplest choice of neural network and retains the Markovian behavior of the FHMM and SLDS, which may be beneficial for interpretability. In some cases, the Markovian assumption is not appropriate, for example if data exhibit strong stereotyped temporal patterns. In such cases, we use RNNs.
We then optimize the ELBO, following Kingma and Welling, (2014). We treat $\sigma_{p}$, $\{\theta_{k}\}$ and $\{A^{(k)}\}$ as parameters to be optimized using stochastic gradient descent. To encourage sparsity in the inferred latent structure, we add an L1 penalty $\lambda\sum_{t}\sum_{k}h_{t}^{(k)}$ on the learned $h_{t}^{(k)}$ to the ELBO. The per-epoch computational cost of our model scales as O(TKD).
@@ 4 Experimental evaluation
In this section, we evaluate FSLDS on a series of real and synthetic datasets. Unless otherwise stated, we use an MLP to model the functions $f_{p}$, $f_{q}$ and $f_{z}$. We found that, in the absence of clear stereotypy, there was no clear advantage with using a more flexible model such as an RNN. Complete implementation details are found in Appendix A of the Supplementary Material (Gong et al.,, 2025). For experiments on real data, we conducted posterior checks using Pareto Smoothed Importance Sampling (PSIS, Yao et al.,, 2018). In all cases we met the proposed stability check ($\hat{k}<0.7$.)
@@
(a) Latent features $\theta_{1:4}$.
@@
(e) Combined rate $\sum_{k=1}^{4}\theta_{k}h_{t}^{(k)}e^{z_{t}^{(k)}}$.
@@ 4.1 Simulated Data
We begin by exploring the performance of FSLDS in a synthetic example where the ground truth features are known. This allows us to explore the ability to recover latent structure and showcase the benefits of using a factorial model over a simpler SLDS. We generate $T=1000$ time steps, with $K=4$ latent, 16-dimensional features $\theta_{k}$, as shown in Figure 3(a). Each figure has a latent continuous trajectory $z_{t}^{(k)}$ drawn either from a straight line or a sinusoid, as shown in Figure 3(c). These were combined with manually generated binary trajectories $h_{t}^{(k)}$(Figure 3(b), Figure 3(d)) to obtain the rate shown in Figure 3(e). This rate was used to sample random Poisson observations.
In Figure 4, we show the posterior means obtained via the inference procedure described in Section 3.3, using a regularization coefficient $\lambda=0.1$ (chosen via cross-validation as described in Appendix A of the Supplementary Material (Gong et al.,, 2025). The results indicate that the Poisson rates are accurately recovered and are appropriately decomposed into the true generating features, demonstrating the model’s ability to capture the underlying structure of the data. Even though we set $K=6$, the model correctly identifies the four latent features shown in Figure 3. We note that, while the RFSM uses a continuous-valued indicator $h_{t}^{(k)}$, in practice the values obtained are very close to zero or one (Figure 4(b)); we see similar behavior in all other experiments in this section.
@@
Figure 5: Latent states obtained using rSLDS. Left hand plot shows the ground truth latent states (corresponding to the thirteen unique combinations of four additive features shown in Figure 3(b)). Remaining plots show the results with varying numbers of latent states, from four to sixteen.
To attempt to learn this true, non-additive structure, we run recurrent SLDS (rSLDS), using the ssm package associated with the original paper (Linderman et al.,, 2017). We test rSLDS with $K\in[4,16]$ states (with the correct number being 13) and 4 latent features, taking the best ELBO out of ten random seeds. The resulting state assignments are shown in Figure 5. We see that we cannot recover the ground truth. We find that rSLDS tends to miss low-occupancy features and combine multiple features. Even though most of the plots show between two and four main features, these features are not consistent between plots. The results also tend to be noisier than those obtained using FSLDS, particularly as the number of features increases.
In the top row of Table 1, we compare the average one-step-ahead predictive RMSE of FSLDS, rSLDS, and a linear dynamical system (LDS). Each model was trained on the first 700 observations, and evaluated on the remaining 300. Both rSLDS and LDS are implemented using the ssm package (Linderman et al.,, 2017). For rSLDS, we used $K=14$ latent states (selected via cross-validation) and 4 latent features (to match the ground truth). We see that, in addition to yielding more interpretable structure as described previously, FSLDS has a notably lower held-out RMSE than the rSLDS and LDS.
Dataset FSLDS rSLDS LDS
Simulation
Simulation 2.4766 5.4494 5.1786
Hippocampal Data
TC043 DIV14 9.6320 10.9240 11.5870
TC165 DIV14 8.2190 9.2357 11.6800
iPSC Data
P1B2 3.8150 3.5248 4.9054
P5A2 4.9150 6.4342 2.9484
P3B2 3.7388 2.2556 2.5836
Table 1: RMSE of held-out observations, for simulated data (top); murine hippocampal data (middle) and human iPSC data (bottom).
@@ 4.2 Evaluation of microelectrode array spike train data
To evaluate the performance of our algorithm on real-world neuronal connectivity data, we look at two datasets obtained via MEA recordings. For both datasets, we use preprocessed data from  Sit et al., (2024). To obtain this data, spikes (action potentials) were detected from the raw MEA voltage time series, based on the action potential waveform using the continuous wavelet transform method implemented in MEA-NAP, a MATLAB pipeline for MEA network analysis (Sit et al.,, 2024). [^3] https://github.com/SAND-Lab/MEA-NAP The voltage timeseries were acquired at a sampling rate of 12500 Hz (human iPSC-dervied neuronal cultures) and 25000 Hz (murine hippocampal cultures). Action potentials are approximatedly 1ms in duration. For our analysis, we aggregated the spike train data to 1 second time bins.
@@ 4.2.1 Exploring properties of the inferred structure using murine hippocampal recordings
(d) Combined rate $\sum_{k}\theta_{k}h_{t}^{(k)}e^{z_{t}^{(k)}}$.
(e) Latent subgraphs $\theta_{k}$.
@@
Figure 6: Murine hippocampal data (recording TC043 DIV14). Figure 6(a) shows the spike counts (color bar, truncated to the 95th percentile value) from a 10-minute MEA recording of a days-in-vitro (DIV) 14 primary hippocampal culture. The activity observed in each electrode (columns) is shown over time (y-axis unit is seconds). Spike count binned by 1 second intervals. Figure 6(b) shows the binary feature indicators $h_{t}^{(k)}$, indicating which of the latent features are active at each time point. Figure 6(c) shows the expression level of each of the latent features $\theta_{k}$. Figure 6(d) shows the inferred firing rate. Figure 6(e) shows the relative contribution of each electrode, in the spatial arrangement of the MEA, for a subset of the latent states $\theta_{k}$, including the background (subnetwork 0) and most active subnetworks (1-4). The relative contribution (node color) is scaled for each subnetwork. The predictive RMSE of the last 50 seconds is 9.632.
We first consider qualitative analyses based on MEA recordings from two 2D murine hippocampal cultures (TC043 DIV14 and TC165 DIV14), first published in (Schroeter et al.,, 2015), using the spike train data extracted by Sit et al., (2024). Results for TC165 are included in Appendix C.1 of the Supplementary Material (Gong et al.,, 2025). We visualize the neuronal activity from these two recordings in Figures 6(a) and Figure 1a in the Supplementary Material. Since we see clear stereotypy that would not be captured using a Markovian switching model, we choose to use RNNs to model $f_{p}$, $f_{q}$ and $f_{z}$.  [^4] Fourier analysis suggests TC043 DIV14 and TC165 DIV14 have an average period of 16.9 seconds and 33.5 seconds respectively. We used a third recording, from culture TC050 DIV21, as a validation set to select the regularization parameter $\lambda$ as described in Appendix A in the Supplementary Material (Gong et al.,, 2025), yielding a value of $\lambda=0.1$.
@@
The remaining subplots in Figure 6 (and Figure 1 in the Supplementary Material) show the latent structure obtained using FSLDS. In each case, we see the activity decomposed into distinct subgraphs (Figure 6(e) and Figure 1e in the Supplementary Material). Some features are active throughout most of the recording, but with varying levels of activity (e.g., features 1 and 2 in Figure 6). Others are active in short bursts, such as features 5 and 6 in Figure 6. Posterior predictive checks using PSIS yielded $\hat{k}=0.51$ for TC043 DIV14 and $\hat{k}=0.34$ for TC165 DIV14, suggesting approximate convergence.
We next look at quantitative performance of FSLDS. In the middle row of Table 1, we calculate average one-step-ahead predictive RMSE for the two cultures described above, using FSLDS, rSLDS, and LDS. In each case, we trained on the first 550 observations, and evaluated on the last 50. For rSLDS we select the number of latent states and latent features using held-out cross-validation on culture TC050 DIV21. In both cases, we see better predictive performance using FSLDS, indicating that we do not sacrifice predictive performance in favor of interpretability.
One concern might be that the number of inferred latent features will grow with sequence length regardless of the true latent dimensionality. In Appendix B in the Supplementary Material (Gong et al.,, 2025), we demonstrate that this is not the case: We show that the structure uncovered in Figure 6 (and Figure 1 in the Supplementary Material) persists if we vary the recording length.
Figure 7: Number of activated features across weeks in culture. Violin plots with scatter plots of number of active features for each culture (circles or stars, n=15 cultures) at four developmental time points days-in-vitro (DIV) 14, 21, 28, and 35. Lines show developmental trajectory of number of latent features within individual cultures. Three examples are highlighted with stars and thicker colored lines: P1B2 (pink, see also Figure 9), P5A2 (orange, see also Figure 4 in the Supplementary Material (Gong et al.,, 2025)
), and P3B1 (purple, see also Figure 5 in the Supplementary Material).
@@ 4.2.2 Exploring evolution of connectivity over longer time periods
Next, we consider a dataset of MEA recordings from human induced pluripotent stem cell (iPSC)-derived neurogenin-2 (NGN2) cortical neurons cultured on 64-electrode MEAs (n=15 cultures), first analyzed in Sit et al., (2024). Each culture was recorded for a period of ten minutes once per week, over a period of four weeks, at days-in-vitro (DIV)14, 21, 28 and 35 for $n=15$ cortical cultures. Spike counts were binned into one-second intervals. To investigate subnetwork development in vitro, we concatenated the data from all four weeks for each culture. We randomly select one dataset, P1A2, as the validation set to choose the penalty $\lambda$, as described in Appendix A in the Supplementary Material (Gong et al.,, 2025) Here $\lambda$ is selected as 0.02, which is associated with the lowest reconstruction error.
@@
In Figure 7, we show the average number of active features (or subnetworks) for each week, for each recording. In almost all cultures, we see an increase in the number of active features across the four weeks, suggesting a developmental process. This coincides with an increase in the number of active electrodes and the number of action potentials detected by these electrodes (Sit et al.,, 2024). However, this pattern diverges at DIV 35, where a bimodal behaviour in the dataset is observed. While a few cultures continue to show a similar number (or increase) in the number of subgraphs at DIV 35 compared to DIV 28 (e.g., P5A2, shown in Figure 5 in Appendix C.2 in the Supplementary Material (Gong et al.,, 2025)), in other cultures, this number decreases (e.g., Figure 9). These likely represent more mature cultures in which the increase in the degree of connectivity has led to a consolidation of subgraphs between DIV 28-35 (see Sit et al.,, 2024)).
@@
(d) Combined rate $\sum_{k}\theta_{k}h_{t}^{(k)}e^{z_{t}^{(k)}}$.
(e) Latent subgraphs $\theta_{k}$.
(f) Mean network activity.
@@
Figure 9: Development of human iPSC-derived NGN2 cortical neurons in culture P1B2. Figure 9(a) shows the raster plot from the concatenated spike train data from days-in-vitro (DIV) 14, 21, 28, and 35 with the spike count (binned in 1 second intervals) for the 64 electrodes (columns) over the four 10-minute recordings (x-axis unit is seconds). Figure 9(b) shows the latent features (columns) for the 10-minute MEA recordings from the four developmental time points (rows). Color bar represents the probability of the feature being on (yellow) or off (purple). Figure 9(c) shows the combined activity (color bar) revealing the temporal dynamics within the on and off switching of the latent states within each MEA recording. Lighter colors (pink to yellow) show dominant features in contrast to features that are off (dark blue). Figure 9(d) shows the inferred firing rate (color bar). Figure 9(e) shows heatmaps of the relative contribution (color) of individual electrodes (circles) to the subgraphs in the spatial arrangement of the MEA for subnetworks 0-7. Figure 9(f) shows heatmaps in the spatial arrangement of the MEA of the mean spike count (per second) per electrode (color) at the four developmental time points. PSIS coefficient is -0.09<0.7.
@@
In the first culture, P1B2 (Figure 9), we see an initial increase in the number of subnetworks from DIV 14-28, followed by a mild consolidation of the network structure into new activity patterns at DIV 35. This illustrates the ability of the model to detect biologically relevant changes in subnetworks as the neuronal networks mature. The model detects seven different features (Figure 9(b)9(c)), and the latent subgraphs reveal different combinations of electrodes detecting activity in these subnetworks (Figure 9(e)). From DIV 14 to DIV 28, we observe an increasing number of different features that are turned on (from 1 to 3). In addition, we see one feature introduced at DIV 21 remains active during DIV 28. However, at DIV 35, one completely different feature is activated. These findings are consistent with an increase in network activity (Figure 9(f)) and density between DIV14-28, illustrated in Sit et al., (2024), and an increase in the mean nodal participation coefficient between DIV 28-35. When considering the spatial arrangement of the subnetworks (Figure 9(e)), we also observe a shift in participating nodes. During the first three weeks, the active nodes are more likely to be located in the middle or right side of the MEA, while at DIV35, the activity has shifted to the left. This may indicate that some subnetworks may increase or decrease in dominance in the overall network activity as the cultures mature.
@@
Figures for two additional cultures are shown in Appendix C.2 in the Supplementary Material (Gong et al.,, 2025) The second culture, P5A2 (Figure 4 in the Supplementary Material), clearly demonstrates an increase in network complexity from DIV 14-35 (Figures 4b–4c in the Supplementary Material), without the consolidation behavior seen in P1B2 at DIV 35. With no active features observed at DIV14, subnetwork 0 (Figure 4e in the Supplementary Material) mainly comes from the baseline noise. As the neuronal networks mature, more features are activated, forming more complex subnetworks (Figures 4c–4e in the Supplementary Material). We hypothesize that P5A2 may be following a similar developmental trajectory to P1B2 , but with a slower start (as evidenced by the lack of correlated activity at DIV 14). As with P1B2 , we observe the overall network structure becoming more complex as additional subnetworks are activated.